Data for the decomposition of dinitrogen monoxide on a gold surface at 1173 K are given below.

N2O --> N2 + ½ O2

Time (min)

15.0

30.0

80.0

120.0

[N2O] (mol/L)

0.0835

0.0680

0.0350

0.0220

Using a graphing calculator, plot a graph of 1/[N 2O] versus time. What is the linear regression correlation constant, r? Make sure to include at least 3 significant digits.(1)


1
Expert's answer
2021-01-15T06:46:51-0500


The linear regression correlation constant is

Σ X = 250 , Σ Y = 0.2085 ,Σ XY = 8.7325 ,

Σ X2 = 21925 , Σ Y2 = 0.01330525


r = "\frac{n•ΣXY - ΣX•ΣY}{\sqrt{[nΣX^2-(ΣY^2)]•[nΣY^2-(ΣY^2)]}}"


r = "\frac{4×8.7325-(245×0.2025)}{\sqrt{[4×21924-245^2]•[4×0.01330525-0.2085^2]}}"


r= -0.9834